A light hangs in 15 feet directly above a straight walk on which a man is 6 feet tall is walking. How fast is the end of the man's shadow travelling when he is walking away from the light at the rate of 3 miles per hour?

1
Expert's answer
2022-02-19T07:02:15-0500

Let |DF|=h – man height, |BC|=H – light height, V=3 miles/hour – walking rate, |AF|=S – shadow size.

|DE|=V"\times" t – horizontal distance from light on time t.


Triangles ADF and DBE are similar.

So "\frac{S}{h}"="\frac{|DE|}{H-h}" /(H-h) 


"\frac{S}{h}"="\frac{(V\times t)}{(H-h)}"   


S= "\frac{(V\times t\times h)}{(H-h)}"


From last expression or differentiating this expression by t shadow size rate Sh = "\frac{(V\times H)}{(H-h)}"

="\frac{(3\times6)}{(15-6)}" = 2 miles/hour.   


The rate of the end of the man's shadow is the velocity of point A


Sa = "\frac{(V\times H)}{(H-h)}" +V = "\frac{(V\times H)}{(H-h)}" = "\frac{(3\times5)}{(15-5)}" (15-6) = 5 miles/hour.


Answer


The rate of the end of the man's shadow Sa = 5 miles/hour.